OpenStudy (anonymous):

A television normally sells for $320. This week, it's on sale at 60% of the normal price. What is the sale price? A) $128 B) $192 C) $256 D) $301

OpenStudy (anonymous):

@confluxepic

OpenStudy (calculusxy):

$320 x 0.6 = x $320 - x = your answer

OpenStudy (confluxepic):

Of means times.

OpenStudy (confluxepic):

60% of 320 would be 60% times 320.

563blackghost (563blackghost):

$320 x .60 = tax $320 - tax = discount total

OpenStudy (confluxepic):

@HappySprout

OpenStudy (anonymous):

so the answer is...

OpenStudy (anonymous):

Thank you. Well... It is A

OpenStudy (confluxepic):

No.

OpenStudy (confluxepic):

\[.60 \times 320=?\]

OpenStudy (anonymous):

192

OpenStudy (anonymous):

no, shes right. I REALLY hope you're a she. that would be SUPER awkward.

OpenStudy (anonymous):

I am a she. lol!

OpenStudy (calculusxy):

Yeah @HappySprout is correct.

OpenStudy (anonymous):

I agree

OpenStudy (anonymous):

you have to take away the discount from the original price.

OpenStudy (anonymous):

are you in k12?

OpenStudy (confluxepic):

That's correct.

OpenStudy (confluxepic):

\[.60 \times 320=192\]

OpenStudy (anonymous):

So is 192 correct?

OpenStudy (anonymous):

I am so confused.

OpenStudy (anonymous):

@confluxepic Is it right or wrong

OpenStudy (confluxepic):

\(\color{blue}{\text{Originally Posted by}}\) @confluxepic That's correct. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (anonymous):

Ok thank you!

OpenStudy (anonymous):

ok i have more. I'll tag you.